ganeshie8 (ganeshie8):

show that \[\large \int\limits_0^{\infty} \dfrac{x^{a}\,dx}{1+x^b} = \frac{\pi}{b\sin(\pi(a+1)/b )}\] \(a\ge 0\) and \(b\gt a+1\)

OpenStudy (anonymous):

what have u played with beta :O

ganeshie8 (ganeshie8):

kindof.. its going to be simple with a clever substitution

ganeshie8 (ganeshie8):

btw that integral only converges for b>a+1 by comparison test so assume thats the case

OpenStudy (anonymous):

\((1+x)^b=1+x^b+\sum _{k=1}^{b-1}x^{k}\\ 1+x^b=\sum _{k=1}^{b-1}x^{k}-(1+x)^b \)

ganeshie8 (ganeshie8):

Interesting, but what good is that

OpenStudy (anonymous):

\(\large \large \int\limits_0^{\infty} \dfrac{x^{a}\,dx}{1+x^b} = \large \large \int\limits_0^{\infty} \dfrac{x^{a}\,dx}{\sum _{k=1}^{b-1}x^{k}-(1+x)^b } \) by partial fraction \(\large \large \dfrac{x^{a}}{1+x^b} = \large \large \dfrac{A }{\sum _{k=1}^{b-1}x^{k} } - \dfrac{B }{(1+x)^b } \) i dont even think if this is possible in this way but i'll give a try

ganeshie8 (ganeshie8):

You're assuming a and b are integers

OpenStudy (anonymous):

eh , never mind then :\

ganeshie8 (ganeshie8):

its okay lets assume they are integers

OpenStudy (anonymous):

i dont know how to do the fraction thingy so i'll try something else

OpenStudy (anonymous):

i got something nice

ganeshie8 (ganeshie8):

okay im waiting :)

OpenStudy (anonymous):

ugh nvm i made mistake somewhere leave it

ganeshie8 (ganeshie8):

Ohk.. I'll leave the question open and check in the morning, cya!

OpenStudy (xapproachesinfinity):

just say we don't need any of those Beta and Gamma functions here

OpenStudy (anonymous):

\(\large \text{assume :-} \\ u=x^b , \text{ then } x=u^{\frac{1}{b}}\text{ and } \large \,dx=\frac{1}{b}\frac{du}{x^{b-1}}\\ \\ \large \begin{align*} \int _0^\infty \frac{x^a}{1+x^b} &= \frac{1}{b} \int _0^\infty \frac{x^a}{(1+u)x^{b-1}}\,du\\ &=\frac{1}{b} \int _0^\infty \frac{x^{a-b+1}}{(1+u) }\,du \\ &= \frac{1}{b} \int _0^\infty \frac{u^{\frac{a +1}{b}-1}}{(1+u) }\,du \end{align*} \\ \large \text{ let } m =\frac{a+1}{b} \text{ and } n+m=1 \text{ then }n=1- \frac{a+1}{b} \\\large \)

OpenStudy (anonymous):

\(\begin{align*} \frac{1}{b} \int _0^\infty \frac{u^{\frac{a +1}{b}-1}}{(1+u) }\,du &= \frac{1}{b} \int _0^\infty \frac{u^{m-1} }{ (1+u)^{n+m } }\,du \\ &=\frac{1}{b} \beta(m,n) \\ &=\frac{1}{b} \beta(\frac{a+1}{b},1-\frac{a+1}{b}) \\ &= \frac{1}{b} \frac{\Gamma(\frac{a+1}{b})\Gamma( 1-\frac{a+1}{b})}{\Gamma (\frac{a+1}{b}+1-\frac{a+1}{b} )}\\ &=\frac{1}{b} \frac{\Gamma(\frac{a+1}{b})\Gamma( 1-\frac{a+1}{b})}{\Gamma (1 )}\\ &=\frac{1}{b} \frac{\pi}{\sin(\pi (\frac{a+1}{b}))} \end{align*} \)

OpenStudy (anonymous):

yaaaaaaaaaaaaaaaay :3 done

OpenStudy (anonymous):

note that \(\Gamma(1)=1\)

OpenStudy (anonymous):

and used Euler's reflection formula for last line

OpenStudy (michele_laino):

we have to go in the complex plane,namely I set: \[x ^{(b/2)}=z\] so your integral becomes: \[\frac{ 2 }{ b } \int\limits_{0}^{\infty}\frac{ z ^{2a/b+2/b-1} }{ 1+z ^{2} }dz\] now that integral is equal to: \[\frac{ 2\pi }{ 2 }(Resf(i)+Resf(-i))\] where Res is the residual value of f(z) and f(z) is our integrand above. after a standard calculation, we get: \[\frac{ \pi b }{ 4 }\left( e ^{i \pi(a+1)/b}-e ^{-i \pi (a+1)/b} \right)=\] \[=\frac{ \pi b }{4 }*2i \sin \left( \pi \frac{ a+1 }{ b } \right)\] and taking the real part we can write: \[\frac{ \pi b }{ 2 }*\sin \left( \pi \frac{ a+1 }{ b } \right)\]

OpenStudy (michele_laino):

please, note that I got sin at numerator, not at denominator

OpenStudy (michele_laino):

is it possible that?

OpenStudy (anonymous):

wait u have little problem with residues , did u use Taylor series to get them ?

OpenStudy (michele_laino):

no, only the theory of integration of complex functions

OpenStudy (michele_laino):

oops.. sorry the coefficient is: \[\frac{ \pi }{ b }\] not \[\frac{ \pi b }{ 2 }\]

ganeshie8 (ganeshie8):

Hahahaa @Marki gave u both beta and gamma functions @xapproachesinfinity :P

OpenStudy (anonymous):

:|